(1 point) Find the average value of the function g(x) = 10:02 V1+? on the interval [0,2]. gave


Find the average value of the function g(x)=10x2 \sqrt{1+x3} on the interval [0,2].


gave=
Answer #1

8(x) = 10x² 1 + x The average value of the function is 2 1 = Save (10x V1+xdx 2 0 10 = | V1+x)(xdx) let, 1+ x = u 3x*dx = du